# Opponent Shaping for Antibody Development

### v6

@@ 2.3.2 Amino Acid Distribution in Shapers and Myopic Antibodies
To further understand the performance differences between myopic antibodies and shapers, we analyse how the amino acid distributions of the antibodies change with the optimisation horizon $H$. Within our model, each amino acid is solely characterised by its binding strength to other amino acids as defined by the Miyazawa-Jernigan energy potential matrix [14]. However, despite this simplicity, we still see interesting patterns in the amino acid distribution. Figure 4 showcases the results of our experiment.
@@
Antibodies optimised with longer horizons, especially the $H=100$ shapers exhibit a more uniform distribution of amino acids, while those with shorter horizons show a tendency to cluster around amino acids associated with either high or low binding energies. The flatter distribution of long-horizon shapers suggests a more diverse and balanced approach to viral antigen binding. We hypothesise that this strategy helps to preserve robustness against viral mutations. By maintaining a more even distribution across energy levels, these antibodies may be less susceptible to viral escape.
@@
Figure 4: Distribution of amino acids in myopic antibodies and shapers. The antibodies are optimised for $N=30$ steps using different viral escape horizons $H$. Longer horizon shapers push the amino acid distribution closer to a uniform distribution.
2.3.3 Influence of Antibody Shapers on Binding Poses
@@
Figure 5: Influence of antibody shapers on binding poses. a Average pose matrices between antibodies optimised using different horizons and the virus at various stages of its escape. The escape steps increase from left to right and the horizon increases from top to bottom. The full grid of matrices with more antibody horizons and virus escape steps is available in the Supplementary Information, Figure 7. b, c Aggregated sum of pose matrices w.r.t the antibody axis (b) and w.r.t the virus axis (c). The plots show a change in the interaction counts in the poses from the viral escape step 00 to $100$. Red indicates a decrease in the interaction count and green an increase.
@@
In the Absolut! framework [11], binding poses are defined as sets of interacting residue pairs between the antibody and the antigen. The binding energy of a pose is calculated by populating these residue locations with the amino acid sequences of both the antibody and the virus and then summing the pairwise interaction energies defined by [14]. Absolut! considers a vast number of possible poses (on the order of $10^{6}$) and determines the overall interaction energy as the energy of the minimum pose, refer to Methods Section 4.4 for more details. Importantly, only a small part of the antigen sequence contributes to this minimum energy pose.
@@
As both the virus and the antibody mutate during our optimisation process, the lowest energy pose can change. To capture these dynamics, we introduced the concept of a pose matrix: a $20\times 20$ matrix22220202020 is the number of possible amino acids. where one dimension corresponds to the antibody amino acids and the other to the viral amino acids, both contributing to the lowest energy pose. The entries in this matrix represent the number of interactions between specific amino acid pairs.
@@
Figure 5a presents average pose matrices from multiple optimisation runs of both myopic antibodies and long-horizon shapers. We observe two key trends. First, as viral escape steps increase (top row vs bottom row), the pose matrices become more “diffused”. This is expected, as the virus explores more “pose possibilities” through mutations during escape. Second, as the horizon of antibody optimisation increases, the poses also become more “diffused”. This is particularly interesting, as all antibodies have the same number of mutations regardless of the horizon, suggesting that this diffusion might relate to the increased robustness of shapers.
@@
To further understand these pose dynamics, we aggregate the pose matrices along the antibody axis (Figure 5b) and the virus axis (Figure 5c). These figures show the change in interaction counts between viral escape steps 00 and $100$. Figure 5b shows that as the virus escapes it includes more of the antibody’s lowest binding amino acids (particularly K, Lysine) in the pose. Notably, long-horizon shapers, especially $H=100$ shapers, are most effective at preventing this increase in K (Lysine) interactions. Furthermore, Figure 5c shows another viral escape strategy, where the virus removes its high-binding amino acids I (Isoleucine) and M (Methionine) from the pose. Again, $H=100$ shapers are most successful in counteracting this trend, although they cannot completely prevent it.
@@
3 Discussion
This study introduces a novel framework for designing antibodies using the principles of opponent shaping [9] that can effectively combat evolving pathogens. The core of our framework is a two-player game that models the interaction between antibodies and viruses. Applying opponent shaping to the game allows us to optimise for antibody shapers that anticipate and influence viral evolution.
Our proposed framework is adaptable. While we demonstrated its efficacy using the Absolut! [11] binding simulator accelerated with JAX [13] for payoff estimation, and our model of the simulated viral escape with an evolutionary process, these components are interchangeable. As more accurate and efficient binding simulators or viral evolution models become available, they can be integrated into our framework, potentially enhancing its biological applicability.
Our results demonstrate a clear advantage of antibody shapers over myopic antibodies when considering long horizons of viral escape. We observed that myopic antibodies, while initially effective, struggle to maintain their efficacy as the virus evolves. In contrast, shaper antibodies, particularly those optimized for longer horizons, show sustained performance over extended periods of viral evolution, see Figure 2. A notable computational trade-off emerges in the performance of antibody shapers. While longer optimisation horizons generally lead to better long-term performance, they come at an increased computational cost. Figure 2d offers guidance for selecting optimal optimisation horizons based on available computational resources and runtime of a binding simulation which is particularly valuable for future applications of our framework.
Furthermore, our analysis reveals that long-horizon shapers shape the virus into more exploitable variants as opposed to just becoming robust to any virus, see Figure 3. In fact, the short-horizon shapers bind stronger to the viruses induced by the long-horizon shapers than the long-horizon shapers themselves. This suggests exploring a mixture of antibodies, some optimised for virus shaping and others for binding, leveraging the benefits of both shaping the viral evolutionary landscape and maintaining strong binding efficacy. Interestingly we observe a similar “strategy” deployed by the human body: a longitudinal study on rare patients who naturally developed broadly neutralizing antibodies (bnAbs) to HIV showed that early neutralizing antibodies induce HIV escape to those antibodies, resulting in enhanced binding to other antibodies [23]. In the future, our approach could enable us to better understand and develop therapies for similar settings by simulating a game with more than one antibody playing against the virus, drawing inspiration from broader research on coevolving systems and their complex interactions [24, 25].
To better understand the mechanisms underlying the effectiveness of antibody shapers, we conduct an explainability analysis focused on the amino acid distribution in shapers (Figure 4) and on binding poses (Figure 5). While these results offer insights into the strategies employed by our antibody shapers, it’s important to note that they are specific to the Absolut! binding model [11, 14], and the transferability of our findings to more realistic binding simulators and viral escape models remains an open question. Antibodies designed using more accurate binding simulators might be able to employ other strategies in their interactions with viruses, and the strategies we explore in this model may not apply in a more realistic simulation. Nevertheless, the fact that we can explain the behaviour of our antibody shapers is encouraging. It suggests that, even as we move towards more complex and realistic models, we may be able to maintain a degree of interpretability in our approach.
A significant limitation of our current approach is the static structure of the viral antigen. In reality, mutations to the virus induce changes in the structure of the viral antigen. However, in our experiments, we consistently use the original structure of the 2R29 dengue viral antigen [22] and a fixed set of Absolut!-generated docking poses based on that structure. As we introduce mutations to both the virus and the antibody, the identity of the lowest-energy pose changes, but the underlying set of possible poses remains constant. This simplification potentially overlooks important structural dynamics that influence antibody-antigen interactions. Ideally, we could model the structural changes that occur with each introduced mutation to capture a more realistic representation of the evolving virus-antibody interaction landscape. AlphaFold3 has demonstrated the capability to predict structures of antibody-antigen complexes from their sequences with high accuracy [26]. Incorporating methods like AlphaFold3 into our framework would allow for the dynamic updating of antigen structures in response to our simulated mutations. This could enhance the biological relevance of our method.
A consequence of us using the Absolut! framework [11] is that we make certain modelling assumptions. Namely, we focus only on the CDRH3 binding domain of the antibody and consider a fixed length (11 amino acid) region. Regardless of the CDRH3 region being the most critical for both antibody specificity and affinity [15], the other five binding regions (CDR1-3 on both the heavy and light chains) also partially contribute to the binding profiles, either by making direct contacts with the antigen or putting constraints on the conformational space. In future work, we want to extend the binding simulation to all the CDR regions with both light and heavy chains.
Another promising avenue for enhancing our framework lies in the integration of more accurate viral escape forecasting methods [27, 28]. EVEscape [27] successfully anticipated the emergence of COVID-19 variants using only pre-pandemic data. Incorporating such sophisticated viral escape forecasting as the simulated viral escape component of our framework could significantly boost its real-world applicability. This integration can help to bridge the gap between our computational approach and practical therapeutic development.
The implications of our work extend beyond the specific context of antiviral therapies. While our experiments focus on the dengue virus, our framework is adaptable to any evolving pathogen or even the heterogeneity of cancers. This opens up possibilities for applications in various fields where antibody therapies are increasingly important. Several programs aim to develop antibody-based therapies against cancer in both clinical and pre-clinical stages, with a few already being approved and included in the treatments.[29].
As computational models of protein interactions and evolutionary processes continue to improve, the principles of opponent shaping in therapy design are likely to become increasingly relevant. Our work presents a “proof-of-concept” for developing antiviral therapies that not only combat current viral strains but actively shape the evolutionary landscape to our advantage. While significant challenges remain in translating these computational insights into pre-clinical development, our study demonstrates the potential of opponent shaping in designing effective therapies against evolving pathogens.
4 Methods
Our method applies the principles of opponent shaping [9] to a two-player game between the antibody shaper player or agent and a naive virus agent (Figure 1a). In this section, we describe our method in detail. Firstly, we introduce the virus-antibody game in Section 4.1 where we define the action spaces and the payoffs of both players. Secondly, we define the simulated viral evolution process in Section 4.2 where the virus evolves to escape in response to the current antibody. Finally, in Section 4.3 we define the antibody optimisation process where we optimise the antibody shapers in a way that accounts for future viral mutations and learns to influence viral evolution away from escape. Importantly, implementing the virus-antibody game requires a way to determine the strength of binding of any antibody-antigen pair, or as we refer to it later in this section a binding function. In Section 4.4 we describe our binding function choice, based on the Absolut! framework [11].
4.1 Virus-Antibody Game
We frame our setting as a two-player zero-sum game between the virus ($v$) and the antibody ($a$), meaning that the gain of one player is equal to the loss of the other player, and vice-versa. The action space of both agents is their respective amino acid sequence, and the payoffs are defined by Figure 1b and are the negative of each other - i.e. an improvement in antibody payoff results in a corresponding reduction of virus payoff of equal magnitude.
We define the set of 20 amino acids as $\mathbb{A}$. Let $N_{v}$ be the sequence length of the viral antigen, and $N_{a}$ be the sequence length of the antibody. So an action of the virus will be $v\in\mathbb{A}^{N_{v}}$, and an action of the antibody will be $a\in\mathbb{A}^{N_{a}}$. In our experiments, we use the dengue envelope antigen [22] as the virus, it is composed of 97 amino acids, thus $N_{v}=97$. $N_{a}$ is the sequence length of the antibody’s CDRH3 region which Absolut! [11] assumes to be $11$ amino acids long, so $N_{a}=11$.
Let $B:\mathbb{A}^{N_{v}}\times\mathbb{A}^{N_{a}}\rightarrow\mathbb{R}$ be our binding function, which measures the strength of the binding between the antibody and the viral antigen with increasing values corresponding to stronger binding [^3] This is opposite to binding energies, which are smaller for stronger binding.. Intuitively, the main payoff/reward for the antibody $R_{a}$ should be $R_{a}=B(v,a)$ since it describes the capacity of the antibody to bind the virus, and the corresponding reward for the virus should be $R_{v}=-B(v,a)$.
However, such a choice would lead to a globally optimal strategy for the viral antigen to be completely inert, while the antibody would evolve to be maximally “sticky”. Such results fail to capture the biological realities: viruses must retain their ability to bind to host cell receptors to function, and antibodies need to maintain a certain level of specificity to avoid off-target binding [30]. Altogether, to create a biologically relevant and realistic game setting, we must incorporate additional incentives into our objectives for both the antibody and the virus.
To achieve this, we introduce a binding target $t^{+}_{v}$ for the virus. In the real world, this corresponds to the cell surface receptor protein, that is used by the virus to enter the host cell. To mimic the requirement that antibodies should not become self-reactive to the host proteins and trigger auto-immunity we also introduce a dummy anti-target $t^{-}_{a}$ representing a protein for the self, that is a target to which the antibody should not bind.
This establishes the complete antibody payoff that can be used in the framework of our two-player virus-antibody game:
$R_{a}(v,a)=B(v,a)-B(t^{-}_{a},a)-B(v,t^{+}_{v})$
With $R_{v}(v,a)=-R_{a}(v,a)$. In this framework, the antibody gets a higher payoff not only if it binds the virus with higher affinity, but also if the virus fails to bind its receptor or if the antibody has a lower risk of autoimmunity. We can now define the simulated viral escape via evolution. In the following section, we leverage the viral payoff $R_{v}(v,a)$ as a fitness function to model how viruses might evolve to evade antibody binding. This simulates viral trajectories that viruses could take in response to therapeutic pressure.
4.2 Simulated Viral Escape via Evolution
We model the escape of the virus $\hat{v}$ by simulating the viral evolution from a fixed antibody $a$ over a horizon $H$ number of steps. The simulated viral escape via evolution (see Figure 1a) is defined as follows. Given a starting virus $\hat{v}$, the fixed antibody $a$ induces a distribution $\text{Ev}(\hat{v},a)$ over sequences of viruses $\hat{v}^{1},\hat{v}^{2},\hat{v}^{3},\dots$. We write that $\bm{\hat{v}}=\hat{v}^{0},\hat{v}^{1},\hat{v}^{2},\dots,\hat{v}^{H}$, where $\hat{v}^{0}=\hat{v}$ and $H$ is the horizon length. Then a viral escape trajectory is given by $\bm{\hat{v}}\sim\text{Ev}(\hat{v},a)$. Below we define the process of generating the escape trajectories inductively.
At generation $i$, we have a virus $\hat{v}^{i}$. We then generate a population of viruses $v^{i}_{1},v^{i}_{2}\dots v^{i}_{P}$ by duplicating $\hat{v}^{i}$ $P$ times, then randomly applying mutation such that on average there is one amino acid mutation per viral sequence:
$v^{i}_{k}=\hat{v}^{i}\oplus\textit{Mutation}$
With duplicates in the population being considered distinct, the likelihood of a particular variant increases with the number of duplicates. Furthermore, $\beta$ is a constant which reflects how random the selection process is, with $\beta\rightarrow\infty$ reflecting deterministic max-fitness selection. After $H$ generations, $H$ being our chosen horizon, a full escape trajectory $\bm{\hat{v}}=\hat{v}^{0},\hat{v}^{1},\hat{v}^{2},\dots,\hat{v}^{H}$ has been generated and the simulated viral escape process ends.
4.3 Antibody Optimisation
$F^{H}_{\hat{v}}(a)=\mathbb{E}_{\bm{\hat{{v}}}\sim\text{Ev}(\hat{v},a)}\left[\frac{1}{H+1}\sum^{H}_{i=0}R_{a}(\hat{v}^{i},a)\right]$
As opposed to Equation 2 in Section 2.1.3, which uses a simplified notation, this is a more formal definition of the antibody fitness. Note that if $H=0$ this fitness defaults to a naive antibody payoff that ignores viral escape, i.e. $F^{0}_{\hat{v}}(a)=R_{a}(\hat{v},a)$. We refer to this as the myopic objective.
To optimise antibody shapers, we employ Monte Carlo simulations to estimate the antibody fitness, combined with an evolutionary optimisation algorithm. We refer to this process as the antibody optimisation loop, see Figure 1a. In meta-learning terms, this is the outer loop or the meta-loop, contrasting to the inner loop which is the simulated viral escape via evolution.
Given a starting antibody $\hat{a}$, a starting virus $\hat{v}$ and a viral escape horizon $H$, the antibody optimisation process generates a trajectory of antibodies $\bm{\hat{a}}=\hat{a}^{0},\hat{a}^{1},\hat{a}^{2},\dots,\hat{a}^{N}$, where $N$ is the number of antibody optimisation steps (or meta-steps). In the trajectory, $\hat{a}^{0}=\hat{a}$ is the starting antibody and $\hat{a}_{N}$ is the final optimised antibody. This optimisation could, in principle, start from any antibody, but for simplicity we opt to start from purely random antibodies, meaning $\hat{a}$ is random. In most of our experiments $N=30$.
At the start of antibody optimisation step $i$, we have an antibody $\hat{a}^{i}$. We first generate a population of $P_{a}$ antibodies $a^{i}_{1},a^{i}_{2}\dots a^{i}_{P_{a}}$ by taking both the antibody $\hat{a}^{i}$ and $P_{a}-1$ copies of it, with each copy having a single random mutation to the amino acid sequence of $\hat{a}^{i}$. For our experiments, $P_{a}=40$.
We then sample their fitness values $F^{H}_{\hat{v}}(a^{i}_{j})$ with a fixed number $\eta$ of Monte Carlo roll-outs, i.e. we sample $\eta$ independent viral escape trajectories each with horizon $H$ viral escape steps. We found $\eta=5$ to be sufficient. Finally, we select $\hat{a}^{i+1}$ to be the best-performing antibody:
$\hat{a}^{i+1}=\arg\max_{k}\mathbb{E}\left[F_{\hat{v}}^{H}(a^{i}_{k})\right]$
Once the final optimised antibody $\hat{a}^{N}$ is generated, a full optimisation trajectory is complete, $\bm{\hat{a}}=\hat{a}^{0},\hat{a}^{1},\hat{a}^{2},\dots,\hat{a}^{N}$, and the antibody optimisation process finishes.
4.4 Binding Function
@@
The methodology described so far is independent of the choice of the binding function $B:\mathbb{A}^{N_{v}}\times\mathbb{A}^{N_{a}}\rightarrow\mathbb{R}$. In our work, we rely on the Absolut! framework [11] to implement the binding function and in this section, we mathematically formalise the binding energy calculation that Absolut! uses. For further explanation, readers are recommended to refer to the original Absolut! paper [11].
@@
Where $N_{v}$ and $N_{a}$ are taken to be the set of integers up to $N_{v}$ and $N_{a}$ respectively.
The energy of a complex of a virus $v\in\mathbb{A}^{N_{v}}$ and an antibody $a\in\mathbb{A}^{N_{a}}$, in a given pose $(\phi^{v\times a},\phi^{a\times a})\in\Phi$ is defined by sum of the energies of each adjacent residues. The energy between a residue pair is determined by which two amino acids it contains, given by a symmetric interaction matrix $M:\mathbb{A}\times\mathbb{A}\rightarrow\mathbb{R}$, which is determined experimentally [14].
@@
Absolut! generates $S$ through a two-step process. First, Absolut! discretises a given structure of the virus $v$ (or any antigen) which is taken from the PDB [21]. Second, Absolut! does a brute-force search over possible (discretised) poses for an antibody $a$ to join to the viral structure. The exact details are not necessary for this paper, and again we refer interested readers to the original paper.
@@
To give a more concrete example, for this paper, we use the dengue virus antigen [22]. Absolut! gives $\approx 1.5$ million poses for this structure. Absolut! also comes with $\approx 20$ million real-world antibody sequences. When using the base dengue sequence as the antigen, across the $20$ million binding calculations only $1027$ binding poses are ever the minimum. Furthermore, the relevance of each pose drops exponentially. The most relevant pose accounts for $20\%$ of binding configurations, and by using the top $100$ poses one would get the exact same result for binding in $95\%$ of antibodies. This gives us a way to make the computation $1000$ times faster [^5] In practice, the difference is closer to $10,000$, likely due to the GPU having to move less data. for a negligible accuracy drop for this particular antigen sequence.
@@
However, this leads to more errors as soon as we change the viral antigen sequence. Looking at the particular poses which lead to binding does reveal another way to cut down on the total number of poses: all of the poses contain at least 18 pairs of residues. As the interaction matrix $M$ is strictly negative, having more pairs of residues always makes the binding energy of a pose lower, meaning it is more likely to be where binding occurs. Out of the original 1.5 million poses, only approximately 37 thousand (1 in 40) contain 18 or more pairs of residues. When using only these residues, we see no differences across any of the evolutionary simulations. It is possible that a pose with 17 or less pairs is the dominant one for some antigen $v$ with antibody $a$, but if so then such pairs appear to be extremely rare.
@@
Using these methods of pruning poses gives us two subsets of the original set of poses, a larger one which almost exactly matches performance, and another which sometimes differs, but is much faster to compute. We refer to these as the high-resolution and low-resolution binding simulators respectively. Note that for the low-resolution binding simulator, the more mutations the virus undergoes the less accurate it becomes. Furthermore, we also do binding to the antibody anti-target, $t^{-}_{a}$. To account for this, we compute the relevant poses for this anti-target too.
@@
When running experiments, we always train with the low-resolution binding simulator, then perform “verification” with the high-resolution one and these are the results we report throughout the paper. The reason is twofold. Firstly this enables us to run many more evolution experiments. Secondly, this mimics the real-life process of transferring out of simulation to the real world. By showing we transfer from the low-resolution binding simulation to the slower, high-resolution binding simulation, we demonstrate that our results are not extremely specific to the exact simulation we use and that any result will not disappear as soon as a more accurate simulation is used. We emphasise that Absolut! does not represent an accurate model of antibody binding. It is instead a toy simulation to demonstrate our methodology. For example, we do not expect our framework when used with this simulation model to yield highly effective, superior antibodies for real-world applications.
5 Data Availability
For our experiment we use the envelope protein E from the dengue viral antigen, PDB code 2R29 [22]. The structure data is available in the PDB [21] under https://www.rcsb.org/structure/2R29. The Absolut! [11] binding pose data we use for our JAX [13] implementation is all available under https://shorturl.at/NuUfN.
6 Code Availability
All code we use to implement our method and obtain the results presented here is available in the GitHub repository https://github.com/olakalisz/antibody-shapers.
Appendix A Viral Mutation Distribution
Figure 6: Distribution of selected viral mutations when evolved against different antibody horizons. Due to more neutral mutations, the distribution is closer to uniform. Longer horizon antibodies cause fewer mutations on average.
Appendix B Influence of Antibody Shapers on Binding Poses
Figure 7: Full grid of pose matrices, they represent the average pose interactions between myopic antibodies or antibody shapers optimised with different horizons and viruses at different stages of their escape.

### v8

@@ Appendix B Amino Acid Distribution in Shapers and Myopic Antibodies
To further understand the performance differences between myopic antibodies and shapers, we analyse how the amino acid distributions of the antibodies change with the optimisation horizon $H$. Within our model, each amino acid is solely characterised by its binding strength to other amino acids as defined by the Miyazawa-Jernigan energy potential matrix (Miyazawa & Jernigan, 1999). However, despite this simplicity, we still see interesting patterns in the amino acid distribution. Figure B.1 showcases the results of our experiment.
@@
Antibodies optimised with longer horizons, especially the $H=100$ shapers, exhibit a more uniform distribution of amino acids, while those with shorter horizons show a tendency to cluster around amino acids associated with either high or low binding energies. The flatter distribution of long-horizon shapers suggests a more diverse and balanced approach to viral antigen binding. We hypothesise that this strategy helps to preserve robustness against viral mutations. By maintaining a more even distribution across energy levels, these antibodies may be less susceptible to viral escape.
@@
Figure B.1: Distribution of amino acids in myopic antibodies and shapers. The antibodies are optimised for $N=30$ steps using different viral escape horizons $H$. Longer horizon shapers push the amino acid distribution closer to a uniform distribution.
Appendix C Influence of Antibody Shapers on Binding Poses
@@
Figure C.1: Influence of antibody shapers on binding poses. a Average pose matrices between antibodies optimised using different horizons and the virus at various stages of its escape. The escape steps increase from left to right, and the horizon increases from top to bottom. The full grid of matrices with more antibody horizons and virus escape steps is available in the Supplementary Information, Figure C.2. b, c Aggregated sum of pose matrices w.r.t the antibody axis (b) and w.r.t the virus axis (c). The plots show a change in the interaction counts in the poses from the viral escape step 00 to $100$. Red indicates a decrease in the interaction count and green an increase.
@@
In the Absolut! framework (Robert et al., 2022), binding poses are defined as sets of interacting residue pairs between the antibody and the antigen. The binding energy of a pose is calculated by populating these residue locations with the amino acid sequences of both the antibody and the virus and then summing the pairwise interaction energies defined by (Miyazawa & Jernigan, 1999). Absolut! considers a vast number of possible poses (on the order of $10^{6}$) and determines the overall interaction energy as the energy of the minimum pose, refer to Appendix D for more details. Importantly, only a small part of the viral sequence contributes to this minimum energy pose.
@@
As both the virus and the antibody mutate during our optimisation process, the lowest energy pose can change. To capture these dynamics, we introduced the concept of a pose matrix: a $20\times 20$ matrix with one entry for each possible pair of amino acids. One dimension corresponds to the antibody amino acids, and the other to the viral amino acids. The entries in this matrix represent the number of interactions between the specific amino acid pairs in the lowest energy pose for the binding configuration between an antibody and an antigen.
@@
Figure C.1a presents average pose matrices from multiple optimisation runs of both myopic antibodies and long-horizon shapers. We observe two key trends. First, as viral escape steps increase (top row vs bottom row), the pose matrices become more “diffused”. This is expected, as the virus explores more “pose possibilities” through mutations during escape. Second, as the horizon of antibody optimisation increases, the poses also become more “diffused”. This is particularly interesting, as all antibodies have the same number of mutations regardless of the horizon, suggesting that this diffusion might relate to the increased robustness of shapers.
@@
To further understand these pose dynamics, we aggregate the pose matrices along the antibody axis (Figure C.1b) and the virus axis (Figure C.1c). These figures show the change in interaction counts between viral escape steps 00 and $100$. Figure C.1b shows that as the virus escapes it includes more of the antibody’s lowest binding amino acids (particularly K, Lysine) in the pose. Notably, long-horizon shapers, especially $H=100$ shapers, are most effective at preventing this increase in K (Lysine) interactions. Furthermore, Figure C.1c shows another viral escape strategy, where the virus removes its high-binding amino acids I (Isoleucine) and M (Methionine) from the pose. Again, $H=100$ shapers are most successful in counteracting this trend, although they cannot completely prevent it.
Figure C.2: Full grid of pose matrices. They represent the average pose interactions between myopic antibodies or antibody shapers optimised with different horizons and viruses at different stages of their escape.
@@
Appendix D Binding Function
@@
In general, ADIOS is independent of the choice of the binding function $B:\mathbb{A}^{N_{v}}\times\mathbb{A}^{N_{a}}\rightarrow\mathbb{R}$. In our work, we rely on the Absolut! framework (Robert et al., 2022) to implement the binding function. In this section, we mathematically formalise the binding energy calculation that Absolut! uses. For further explanation, readers are recommended to refer to the original Absolut! paper (Robert et al., 2022).
@@
Where $N_{v}$ and $N_{a}$ are taken to be the set of integers up to $N_{v}$ and $N_{a}$, respectively.
The energy of a complex of a virus $v\in\mathbb{A}^{N_{v}}$ and an antibody $a\in\mathbb{A}^{N_{a}}$, in a given pose $(\phi^{v\times a},\phi^{a\times a})\in\Phi$ is defined by sum of the energies of each adjacent residues. The energy between a residue pair is determined by which two amino acids it contains, given by a symmetric interaction matrix $M:\mathbb{A}\times\mathbb{A}\rightarrow\mathbb{R}$, which is determined experimentally (Miyazawa & Jernigan, 1999).
@@
Absolut! generates $S$ through a two-step process. First, Absolut! discretises a given structure of the virus $v$ (or any antigen) which is taken from the PDB (consortium, 2018). Second, Absolut! does a brute-force search over possible (discretised) poses for an antibody $a$ to join to the viral structure. The exact details are not necessary for this paper, we refer interested readers to the original paper.
@@
To give a more concrete example, for this paper, we use the dengue virus antigen (Lok et al., 2008). Absolut! gives $\approx 1.5$ million poses for this structure. Absolut! also comes with $\approx 20$ million real-world antibody sequences. When using the base dengue sequence as the antigen, across the $20$ million binding calculations only $1027$ binding poses are ever the minimum. Furthermore, the relevance of each pose drops exponentially. The most relevant pose accounts for $20\%$ of binding configurations, and by using the top $100$ poses one would get the exact same result for binding in $95\%$ of antibodies. This gives us a way to make the computation $1000$ times faster [^3] In practice, the difference is closer to $10,000$, likely due to the GPU having to move less data. for a negligible accuracy drop for this particular antigen sequence.
@@
However, this leads to more errors as soon as we change the viral antigen sequence. Looking at the particular poses which lead to binding reveals another way to cut down on the total number of poses: all of the poses contain at least 18 pairs of residues. As the interaction matrix $M$ is strictly negative, having more pairs of residues always makes the binding energy of a pose lower, meaning it is more likely to be where binding occurs. Out of the original 1.5 million poses, only approximately 37 thousand (1 in 40) contain 18 or more pairs of residues. When using only these residues, we see no differences across any of the evolutionary simulations. It is possible that a pose with 17 or less pairs is the dominant one for some antigen $v$ with antibody $a$, but if so, then such pairs appear to be extremely rare.
@@
Using these methods of pruning poses gives us two subsets of the original set of poses, a larger one which almost exactly matches performance, and another which sometimes differs, but is much faster to compute. We refer to these as the high-resolution and low-resolution binding simulators, respectively. Note that for the low-resolution binding simulator, the more mutations the virus undergoes, the less accurate it becomes. Furthermore, we also do binding to the antibody anti-target, $t^{-}_{a}$. To account for this, we compute the relevant poses for this anti-target too.
@@
When running dengue virus experiments, we always train with the low-resolution binding simulator, then perform “verification” with the high-resolution one, and these are the results we report throughout the paper. For the other viruses and the bacterium we run “verification” in a differently initialised instantiation of the low-resolution simulator. The reason is twofold. Firstly this enables us to run many more evolution experiments. Secondly, this mimics the real-life process of transferring out of simulation to the real world. By showing we transfer from the low-resolution binding simulation to the slower, high-resolution binding simulation, we demonstrate that our results are not extremely specific to the exact simulation we use and that any result will not disappear as soon as a more accurate simulation is used. We emphasise that Absolut! does not represent an accurate model of antibody binding. It is instead a toy simulation to demonstrate our methodology. For example, we do not expect our framework when used with this simulation model to yield highly effective, superior antibodies for real-world applications.
Appendix E MDP for the Virus–Antibody Game
We formalise a single interaction round between an antibody and a virus as the finite–horizon Markov Decision Process $\mathcal{M}$:
$\mathcal{M}=\bigl{\langle}\mathcal{S},\,\mathcal{A}^{\mathrm{v}},\,\mathcal{A}^{\mathrm{a}},\,P,\,R,\,\mu\bigr{\rangle}.$
  • State space: $\mathcal{S}=\mathbb{A}^{N_{v}}\times\mathbb{A}^{N_{a}},$ where a state $s=(v,a)$ is the pair of viral ($v$) and antibody ($a$) amino-acid sequences.
  • Action spaces (chosen simultaneously):
    $\mathcal{A}^{\mathrm{v}}(s)=\mathbb{A}^{N_{v}},\qquad\mathcal{A}^{\mathrm{a}}(s)=\mathbb{A}^{N_{a}}.$
  • Transition kernel (episode terminates after one step):
    $P\!\bigl{(}s^{\prime}\mid s,a^{\mathrm{v}},a^{\mathrm{a}}\bigr{)}=\mathbf{1}_{s^{\prime}=s_{\star}},\quad s_{\star}\text{ is terminal.}$.
  • Reward vector $R=(R_{\mathrm{v}},R_{\mathrm{a}})$. Given joint action $\bigl{(}a^{\mathrm{v}},a^{\mathrm{a}}\bigr{)}$ in state $s$,
    $R_{\mathrm{a}}\!\bigl{(}s,a^{\mathrm{v}},a^{\mathrm{a}}\bigr{)}=B\!\bigl{(}a^{\mathrm{v}},a^{\mathrm{a}}\bigr{)}-B\!\bigl{(}t_{a}^{-},a^{\mathrm{a}}\bigr{)}-B\!\bigl{(}a^{\mathrm{v}},t_{v}^{+}\bigr{)},\qquad R_{\mathrm{v}}=-R_{\mathrm{a}}.$
  • Initial-state distribution $\mu$ puts mass on the wild-type virus and the initial antibody candidate: $\mu\bigl{(}v_{0},a_{0}\bigr{)}=1.$
Because this is a single-step MDP, the return equals the immediate reward $R$ and the discount factor is irrelevant.